1). there are 3 up escalators and one elrvator to take shoppers from first to second floor there are 3 down escalator and the same elevator to take shoppers from second floor to first floor. in how many ways can a shoppers go from first floor to the second floor and back?
2). how many 11 digits numbers are possible if the first digits are 9 and 6, the fifth one is 7 and the last two are 2 and 3?
that you for your help

Answer by sudhanshu_kmr(1152)


1) no. of ways to go first floor to second = 4
no. of ways to return from second floor = 4
total no. of ways = 4*4 = 16

2) here, no. of places that should be filled = 6, all others are fixed.
no. of ways = 10^6 = 1000000